OpenStudy (anonymous):

Find the marginal cost, marginal revenue, and marginal profit functions. C(x) = 6x; R(x) = 9x − 0.001x^2. And find all values of x for which the marginal profit is zero.

OpenStudy (anonymous):

We're using derivatives. I got the answer figured out for though nvmd. c(x)=6 R(x)= 9-.002x and p(x)= 3-0.002x and then p(x)=0.

OpenStudy (anonymous):

do one thing 0.001x^2 -3x = 0 x^2 - 3000x = 0 x(x-3000) = 0 therefore x = 0 or, x = 3000 so ur answer is either 0 or, 3000

OpenStudy (anonymous):

so u got the answer or not